OpenStudy (anonymous):

Hi someone can help me to put these over a common denominator: 4/7 and 3/4. please

OpenStudy (triciaal):

7 X 4 =28 4/7 = 16/28 AND 3/4 = 21/28

OpenStudy (anonymous):

Thanks triciaal for your help and God bless you

OpenStudy (anonymous):

easiest way to do these is a/n and b/m become ma/mn and nb/mn, where writing them together means multiplying them.

OpenStudy (anonymous):

thanks douglaswinslowcooper